Shamrock Industries has the following patents on its December 31, 2019, balance sheet.

Patent Item

Initial Cost

Date Acquired

Useful Life at Date Acquired

Patent A $41,820 3/1/16 17 years
Patent B $15,600 7/1/17 10 years
Patent C $16,800 9/1/18 4 years

The following events occurred during the year ended December 31, 2020.
1. Research and development costs of $235,000 were incurred during the year.
2. Patent D was purchased on July 1 for $45,372. This patent has a useful life of 91/2 years.
3. As a result of reduced demands for certain products protected by Patent B, a possible impairment of Patent B’s value may have occurred at December 31, 2020. The controller for Shamrock estimates the expected future cash flows from Patent B will be as follows.

Year

Expected Future Cash Flows

2021 $2,050
2022 2,050
2023 2,050

The proper discount rate to be used for these flows is 8%. (Assume that the cash flows occur at the end of the year.)

Click here to view factor tables

Collapse question part

(a)

Correct answer. Your answer is correct.
Compute the total carrying amount of Shamrock’ patents on its December 31, 2019, balance sheet.
Total carrying amount $enter the Total carrying amount in dollars

Entry field with correct answer

SHOW SOLUTION

SHOW ANSWER

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

LINK TO VIDEO

Attempts: 1 of 3 used

Collapse question part

(b)

Compute the total carrying amount of Shamrock' patents on its December 31, 2020, balance sheet.
Total carrying amount $enter the Total carrying amount in dollars
Answer #1
1 Particulars Patent A Patent B Patent C
Cost (A) $41820 $15600 $16800
Date acquired 01-03-2016 01-07-2017 01-09-2018
Month upto 31st Dec 2019
Year 2016 10 0 0
Year 2017 12 6 0
Year 2018 12 12 4
Year 2019 12 12 12
Total months (B) 46 30 16
Life in months © 204 120 48
Amortization per month (A/C) = D $205 $130 $350
Total Amortization upto Dec 31 2019 (D*B)= E $9430 $3900 $5600
Carrying amount as on Dec 31, 2019 (A-E) = F $32390 $11700 $11200
2 Particulars Patent A Patent B Patent C Patent D
Carrying amount as on Dec 31, 2019 $32390 $11200 $45372
Amortization for the year $2460 $4200 $2388
Carrying amount as on Dec 31, 2020 $29930 $6417 $7000 $42984
Carrying amount of patent B $11700
Expected future cash flow @8% $5283
(2050*2.5771)
Loss recognized $6417